ABC Manufacturing Ltd. manufacture electrical parts. Data for two of the company's customers is as follows:

Customer A Customer B
Revenue at list price $592,500 $102,000
Units Sold 75,000 10,000
Unit list price $7.90 $10.20
Cost of goods per unit $5.10 $6.25
Sales discounts $11,850 $1,020
Customer - specific costs
Order-taking $13,600 $1,500
Product-handling $14,100 $1,350
Delivery $9,400 $1,290

Require

a) Prepare a comparative income statement in gross margin format with one column for each customer; present customer-specific costs as period expenses.
Answer #1
Customer A Customer B
Units Sold A            75,000            10,000
Unit Selling Price B                 7.90              10.20
Sales Revenue C = A*B        5,92,500        1,02,000
Cost of Goods Sold (per unit) D 5.1 6.25
Total Cost of Goods Sold E = A * D        3,82,500            62,500
Gross Margin F = C - E        2,10,000            39,500
Order Cost G            13,600              1,500
Product handling H            14,100              1,350
Delivery I              9,400              1,290
Net Income J = F - G - H - I        1,72,900            35,360
Please do upvote if you found the answer useful.
Feel free to reach in the comment section in case of any clarification or queries.